OpenStudy (anonymous):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work.

OpenStudy (anonymous):

@Hero @dan815

OpenStudy (anonymous):

@Jhannybean @satellite73 @Loser66 @freckles @EclipsedStar @King.Void. @sleepyjess can someone help me please?

OpenStudy (sleepyjess):

Do you have an idea on how to start?

OpenStudy (anonymous):

no not at all

OpenStudy (sleepyjess):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

what you mean?

OpenStudy (sleepyjess):

\(x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (sleepyjess):

That is the quadratic formula

OpenStudy (sleepyjess):

It may look intimidating, but it's not too bad once you know how to work it :)

OpenStudy (anonymous):

Okay how do I work it lol..

OpenStudy (sleepyjess):

A quadratic equation is ALWAYS "Ax^2 + Bx + C = y" To use the quadratic formula, you will want to substitute 0 for y, which has already been done. Can you figure out what A, B, and C are in the quadratic equation you have?

OpenStudy (anonymous):

A= 6x^2 B= 12x C= 5

OpenStudy (sleepyjess):

Close, but we just want 6, 12 and 5

OpenStudy (anonymous):

Ohh lol

OpenStudy (sleepyjess):

So now we just substitute in 6 for a, 12 for b, and 5 for c then work out the equation

OpenStudy (anonymous):

I got down to \[x= \frac{ -12\pm \sqrt{24} }{ 12 }\] now what do I do?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (carolinar7):

Now simplify what you have there

OpenStudy (anonymous):

How?

OpenStudy (carolinar7):

\[x=\frac{ -12-+4\sqrt{6} }{ 12 }\]

OpenStudy (carolinar7):

Can you simplify any further?

OpenStudy (anonymous):

Why 4 sqrt7 instead of sqrt24=4.89....?

OpenStudy (anonymous):

I meant sqrt6 sorry

OpenStudy (carolinar7):

Sorry let me correct myself it is 2sqrt6

OpenStudy (carolinar7):

Is this Alg. 1

OpenStudy (anonymous):

How!? And no its pre-calculus

OpenStudy (carolinar7):

what grade ar you in

OpenStudy (anonymous):

@sleepyjess please help me, I don't understand!!

OpenStudy (carolinar7):

me too im in pre-calc in 9th grade

OpenStudy (sleepyjess):

oh goodness, I forgot about the square root simplifying >_<

OpenStudy (anonymous):

How do I do that?

OpenStudy (sleepyjess):

@carolinar7 simplified it for you

OpenStudy (carolinar7):

what do you mean you simplified it for me

OpenStudy (sleepyjess):

you simplified the square root already

OpenStudy (anonymous):

2 sqrt6?

OpenStudy (carolinar7):

It is not completely simplified though, you can simplify further too

OpenStudy (carolinar7):

\[x=\frac{ -6+-\sqrt{6} }{ 6 }\]

OpenStudy (anonymous):

so where do I go from \[x= \frac{ -12 \pm 2\sqrt{6} }{ 12 }\]

OpenStudy (anonymous):

Ohhh how did you get that?

OpenStudy (carolinar7):

You simplify it and divide by 2 because that is what they share in common

OpenStudy (anonymous):

Ohhh okay soo now I do -6/6?

OpenStudy (carolinar7):